# Sheffer polynomials and the s-ordering of exponential boson operators

### v1

@@ 6 Examples
The main result (Theorem 5.3) will now be applied, to produce several new $s$-ordered representations of exponential boson operators ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}\in\mathcal{W}[[\lambda]]$, and of $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}\in\nobreak\mathcal{W}$.
According to the theorem, for any $(L,R)$ one needs the EGF of a certain family of two-point Hsu–Shiue polynomials (of Sheffer type), denoted by $\mathcal{T}(t,z;\allowbreak A,B,r,r^{\prime};s)$. The EGF equals $\bar{g}_{T}(z){\rm e}^{t\bar{f}_{T}(z)}$, where $\bar{g}_{T},\bar{f}_{T}]$ is the inverse in the exponential Riordan group $\mathfrak{R}$ of the parametric $s$-dependent element $g_{T},f_{T}]$ defined in eq. (60). One simply substitutes $(A,B,r,r^{\prime})=(e,1,-L,R)$, where $e\mathrel{:=}\allowbreak L+\nobreak R-\nobreak 1$, and sets $(t,z)=({a^{\dagger}}a,\lambda{a^{\dagger}}^{e})$ (which really means $(t,z)=(\alpha^{*}\alpha,\lambda{\alpha^{*}}^{e})$), to obtain the desired $s$-ordered representation.
@@ 6.1 The case $e=0$
This includes the operators ${\rm e}^{\lambda({a^{\dagger}}a)}$, ${\rm e}^{\lambda(a{a^{\dagger}})}$, with $(L,R)=(1,0)$, $(0,1)$. Substituting $(A,B,r,r^{\prime})=\allowbreak(e,1,-L,R)=\allowbreak(0,1,-1,0)$ into (60) gives
$g_{T}(D;\,s)=\frac{2}{2-D-sD},\qquad z=f_{T}(D;\,s)=\ln\left(\frac{2+D-sD}{2-D-sD}\right).$ (81)
@@
$\bar{g}_{T}(z;\,s)=\frac{2}{1+{\rm e}^{z}-s(1-{\rm e}^{z})},\qquad D=\bar{f}_{T}(z;\,s)=\frac{2({\rm e}^{z}-1)}{1+{\rm e}^{z}-s(1-{\rm e}^{z})}.$ (82)
@@
${\rm e}^{\lambda({a^{\dagger}}a)}=\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}{a^{\dagger}}a\right]}\,\mathclose{:}_{s},$ (83)
which is precisely the original Cahill–Glauber $s$-ordering result, eq. (2). The case $(L,R)=(0,1)$ is similar.
@@ 6.2 The case $e=1$
This includes the operators ${\rm e}^{\lambda({a^{\dagger}}^{2}a)}$, ${\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}})}$, ${\rm e}^{\lambda(a{a^{\dagger}}^{2})}$, with $(L,R)=(2,0)$, $(1,1)$, $(0,2)$. Substituting $(A,B)=(e,1)=(1,1)$ into (60) gives
$g_{T}(D;\,s)=\left\{\begin{aligned} &\frac{4+D^{2}-s^{2}D^{2}}{(2-D-sD)^{2}},&\qquad&(L,R)=(2,0);\\ &\frac{4+D^{2}-s^{2}D^{2}}{(2-sD)^{2}-D^{2}},&\qquad&(L,R)=(1,1);\\ &\frac{4+D^{2}-s^{2}D^{2}}{(2+D-sD)^{2}},&\qquad&(L,R)=(0,2);\end{aligned}\right$. (84)
@@
$\bar{g}_{T}(z;\,s)=\left\{\begin{aligned} &\frac{1+s}{1+2sz+z^{2}+(z+s)\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(2,0);\\ &\frac{1}{\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(1,1);\\ &\frac{1-s}{1+2sz+z^{2}-(z+s)\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(0,2);\end{aligned}\right$. (86)
@@
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{2}a)}&=\mathopen{:}\,\frac{1+s}{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}+(\lambda{a^{\dagger}}+s)\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\\ {\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}})}&=\mathopen{:}\,\frac{1}{\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\\ {\rm e}^{\lambda(a{a^{\dagger}}^{2})}&=\mathopen{:}\,\frac{1-s}{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}-(\lambda{a^{\dagger}}+s)\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\end{aligned}$ (88)
@@
$\begin{aligned} \mathcal{E}\mathrel{:=}\exp\left[({a^{\dagger}}a)\,\frac{2\lambda{a^{\dagger}}}{1+s\lambda{a^{\dagger}}+\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\right],\end{aligned}$ (89)
@@
When $s=\nobreak-1$ and $s=\nobreak+1$, each of these three ordered representations reduces to a normal ordered and an anti-normal ordered one, which are the ones provided by Theorem 4.4. (To set $s=\nobreak-1$ in the first or $s=\nobreak+1$ in the third, a limit must be taken.) But, the three Weyl-ordered representations obtained by setting $s=\nobreak 0$ seem to be new to the literature.
From such identities as those of Theorem 6.1, it is possible to derive explicitly $s$-ordered series expansions of $({a^{\dagger}}^{2}a)^{n}$, $({a^{\dagger}}a{a^{\dagger}})^{n}$, $(a{a^{\dagger}}^{2})^{n}$; at least, when s=1,0,+1s=-1,0,+1. For the case of $({a^{\dagger}}a{a^{\dagger}})^{n}$ when $s=\pm 1$, this has already been done: see the Laguerre-polynomial representations in eqs. (55),(56) above, each term in which is explicitly normal, resp. anti-normal ordered. The expansion of $({a^{\dagger}}a{a^{\dagger}})^{n}$ when $s=\nobreak 0$ is of particular interest: it expresses $({a^{\dagger}}a{a^{\dagger}})^{n}$ as the symmetric (Weyl) quantization of a classical function on the complex $\alpha$-plane.
@@
The coefficients in this expansion, unlike those in (55),(56), are rational rather than integer-valued.
@@
Consider the $(L,R)=(1,1)$ case of Corollary 5.4 when $s=\nobreak 0$, or more concretely the $(L,R)=(1,1)$ case of Theorem 5.5, or the second identity in Theorem 6.1 when $s=\nobreak 0$. They all imply that
$({a^{\dagger}}a{a^{\dagger}})^{n}=\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\,\mathbf{R}[\bar{g}_{T}(z;\,0),\bar{f}_{T}(z;\,0)]_{n}(t={a^{\dagger}}a)}\,\mathclose{:}_{W},$ (90)
where $\bar{g}_{T}(z;0)$ comes from the $(L,R)=(1,1)$ case of (86), and $\bar{f}_{T}(z;0)$ comes similarly from (87). That is,
$({a^{\dagger}}a{a^{\dagger}})^{n}=\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\,\mathbf{R}\left[\frac{1}{\sqrt{1+z^{2}}},\frac{2z}{1+\sqrt{1+z^{2}}}\right]_{n}\!(t={a^{\dagger}}a)}\,\mathclose{:}_{W}.$ (91)
@@
$\mathbf{R}_{o}\left[\frac{1}{\sqrt{1+4z^{2}}},\frac{2z}{1+\sqrt{1+4z^{2}}}\right]_{n,k}\!=(-1)^{(n-k)/2}\binom{n}{(n-k)/2}$ (92)
for its matrix elements is known (the binomial coefficient is zero if $n-\nobreak k$ is odd). Here, $\mathbf{R}_{o}$ in distinction to $\mathbf{R}$ signifies an ordinary (non-exponential) Riordan array, meaning that in the definition of $\mathbf{R}_{o}[d(z),h(z)]_{n,k}$ the $n!/k!$ factor in the definition (27) used in this paper is absent. Altering the normalization and restoring the $n!/k!$ factor yields the claimed formula for $({a^{\dagger}}a{a^{\dagger}})^{n}$. ∎
@@ 6.3 The case $e=2$
This includes the operators ${\rm e}^{\lambda({a^{\dagger}}^{3}a)}$, ${\rm e}^{\lambda({a^{\dagger}}^{2}a{a^{\dagger}})}$, ${\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}}^{2})}$, ${\rm e}^{\lambda(a{a^{\dagger}}^{3})}$, with $(L,R)=(3,0)$, $(2,1)$, $(1,2)$, $(0,3)$. As with $e=0,1$, one substitutes $(A,B)=(e,1)=(2,1)$ and $L,R$ into the definitions in eq. (60) of the functions $g_{T}(D;s)$ and $z=f_{T}(D;s)$, and computes $\bar{g}_{T}(z;s),\allowbreak\bar{f}_{T}(z;s)]$ by inverting $g_{T},f_{T}]$ in the exponential Riordan group. The $s$-ordered representation of the operator is then $\bar{g}_{T}(\lambda{a^{\dagger}}^{e};s)\times\allowbreak\exp\left[({a^{\dagger}}a)\bar{f}_{T}(\lambda{a^{\dagger}}^{e};s)\right]$.
But for each of the four operators, the representation is quite complicated. The stumbling block is that according to (60), $z=f_{T}(D;s)$ is given by
$z=f_{T}(D;s)=\frac{8D(2-sD)}{\left[(2-sD)^{2}-D^{2}\right]^{2}},$ (93)
@@
$(1-s^{2})^{2}z\,D^{4}+8s(1-s^{2})z\,D^{3}-8\left[(1-3s^{2})z-s\right]D^{2}-16(1+2sz)\,D+16z=0.$ (94)
The solution $D=D(z;s)$ satisfying $D=1\cdot z^{1}+\nobreak O(z^{2})$ can be expressed in radicals by employing the quartic formula, but owing to its complexity it is not given here. If $s=\pm 1$, the quartic (94) degenerates to a more easily solved quadratic, which can be viewed as the reason why the normal and anti-normal orderings of each of the four operators, which are supplied by Theorem 4.4, are relatively simple. But for general $s$, and even when $s=\nobreak 0$ (the Weyl ordering case), there seems to be no simplification.
Whether it is possible to express $s$-ordered representations of operators ${\rm e}^{{a^{\dagger}}^{L}a{a^{\dagger}}^{R}}$ with $e=L+\nobreak R-\nobreak 1>2$ in terms of radicals, for general or even specialized $s$, is a question requiring further study.

### v2

@@ 6 Examples
The main result (Theorem 5.3) will now be applied, to produce several new $s$-ordered representations of exponential boson operators ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}\in\mathcal{W}[[\lambda]]$, and of $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}\in\penalty 10000\mathcal{W}$.
[…]
@@ 6.1 The case $e=0$
This includes the operators ${\rm e}^{\lambda({a^{\dagger}}a)}$, ${\rm e}^{\lambda(a{a^{\dagger}})}$, with $(L,R)=(1,0)$, $(0,1)$. Substituting $(A,B,r,r^{\prime})=\allowbreak(e,1,-L,R)=\allowbreak(0,1,-1,0)$ into (5.1) gives
$g_{T}(D;\,s)=\frac{2}{2-D-sD},\qquad z=f_{T}(D;\,s)=\ln\left(\frac{2+D-sD}{2-D-sD}\right).$ (6.1)
@@
$\bar{g}_{T}(z;\,s)=\frac{2}{1+{\rm e}^{z}-s(1-{\rm e}^{z})},\qquad D=\bar{f}_{T}(z;\,s)=\frac{2({\rm e}^{z}-1)}{1+{\rm e}^{z}-s(1-{\rm e}^{z})}.$ (6.2)
@@
${\rm e}^{\lambda({a^{\dagger}}a)}=\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}{a^{\dagger}}a\right]}\,\mathclose{:}_{s},$ (6.3)
which is precisely the original Cahill–Glauber $s$-ordering result, eq. (1.2). The functions $g_{T},f_{T}$ given in (6.1) appear in its inverse, eq. (1.3). The case $(L,R)=(0,1)$ is similar to the case $(L,R)=(1,0)$.
@@ 6.2 The case $e=1$
This includes the operators ${\rm e}^{\lambda({a^{\dagger}}^{2}a)}$, ${\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}})}$, ${\rm e}^{\lambda(a{a^{\dagger}}^{2})}$, with $(L,R)=(2,0)$, $(1,1)$, $(0,2)$. Substituting $(A,B)=(e,1)=(1,1)$ into (5.1) gives
$g_{T}(D;\,s)=\left\{\begin{aligned} &\frac{4+D^{2}-s^{2}D^{2}}{(2-D-sD)^{2}},&\qquad&(L,R)=(2,0);\\ &\frac{4+D^{2}-s^{2}D^{2}}{(2-sD)^{2}-D^{2}},&\qquad&(L,R)=(1,1);\\ &\frac{4+D^{2}-s^{2}D^{2}}{(2+D-sD)^{2}},&\qquad&(L,R)=(0,2);\end{aligned}\right$. (6.4)
@@
$\bar{g}_{T}(z;\,s)=\left\{\begin{aligned} &\frac{1+s}{1+2sz+z^{2}+(z+s)\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(2,0);\\ &\frac{1}{\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(1,1);\\ &\frac{1-s}{1+2sz+z^{2}-(z+s)\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(0,2);\end{aligned}\right$. (6.6)
@@
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{2}a)}&=\mathopen{:}\,\frac{1+s}{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}+(\lambda{a^{\dagger}}+s)\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\\ {\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}})}&=\mathopen{:}\,\frac{1}{\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\\ {\rm e}^{\lambda(a{a^{\dagger}}^{2})}&=\mathopen{:}\,\frac{1-s}{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}-(\lambda{a^{\dagger}}+s)\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\end{aligned}$ (6.8)
@@
$\begin{aligned} \mathcal{E}\mathrel{:=}\exp\left[({a^{\dagger}}a)\,\frac{2\lambda{a^{\dagger}}}{1+s\lambda{a^{\dagger}}+\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\right],\end{aligned}$ (6.9)
@@
[…]
@@
The coefficients in this expansion, unlike those in (4.10),(4.11), are rational rather than integer-valued.
@@
[…]
$({a^{\dagger}}a{a^{\dagger}})^{n}=\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\,\mathbf{R}[\bar{g}_{T}(z;\,0),\bar{f}_{T}(z;\,0)]_{n}(t={a^{\dagger}}a)}\,\mathclose{:}_{W},$ (6.10)
where $\bar{g}_{T}(z;0)$ comes from the $(L,R)=(1,1)$ case of (6.6), and $\bar{f}_{T}(z;0)$ comes similarly from (6.7). That is,
$({a^{\dagger}}a{a^{\dagger}})^{n}=\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\,\mathbf{R}\left[\frac{1}{\sqrt{1+z^{2}}},\frac{2z}{1+\sqrt{1+z^{2}}}\right]_{n}\!(t={a^{\dagger}}a)}\,\mathclose{:}_{W}.$ (6.11)
@@
$\mathbf{R}_{o}\left[\frac{1}{\sqrt{1+4z^{2}}},\frac{2z}{1+\sqrt{1+4z^{2}}}\right]_{n,k}\!=(-1)^{(n-k)/2}\binom{n}{(n-k)/2}$ (6.12)
for its matrix elements is known (the binomial coefficient is zero if $n-\penalty 10000k$ is odd). Here, $\mathbf{R}_{o}$ in distinction to $\mathbf{R}$ signifies an ordinary (non-exponential) Riordan array, meaning that in the definition of $\mathbf{R}_{o}[d(z),h(z)]_{n,k}$ the $n!/k!$ factor in the definition (3.3) used in this paper is absent. Altering the normalization and restoring the $n!/k!$ factor yields the claimed formula for $({a^{\dagger}}a{a^{\dagger}})^{n}$. ∎
@@ 6.3 The case $e=2$
This includes the operators ${\rm e}^{\lambda({a^{\dagger}}^{3}a)}$, ${\rm e}^{\lambda({a^{\dagger}}^{2}a{a^{\dagger}})}$, ${\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}}^{2})}$, ${\rm e}^{\lambda(a{a^{\dagger}}^{3})}$, with $(L,R)=(3,0)$, $(2,1)$, $(1,2)$, $(0,3)$. As with $e=0,1$, one substitutes $(A,B)=(e,1)=(2,1)$ and $L,R$ into the definitions in eq. (5.1) of the functions $g_{T}(D;s)$ and $z=f_{T}(D;s)$, and computes $\bar{g}_{T}(z;s),\allowbreak\bar{f}_{T}(z;s)]$ by inverting $g_{T},f_{T}]$ in the exponential Riordan group. The $s$-ordered representation of the operator is then $\bar{g}_{T}(\lambda{a^{\dagger}}^{e};s)\times\allowbreak\exp\left[({a^{\dagger}}a)\bar{f}_{T}(\lambda{a^{\dagger}}^{e};s)\right]$.
But for each of the four operators, the representation is quite complicated. The stumbling block is that according to (5.1), $z=f_{T}(D;s)$ is given by
$z=f_{T}(D;s)=\frac{8D(2-sD)}{\left[(2-sD)^{2}-D^{2}\right]^{2}},$ (6.13)
@@
$(1-s^{2})^{2}z\,D^{4}+8s(1-s^{2})z\,D^{3}-8\left[(1-3s^{2})z-s\right]D^{2}-16(1+2sz)\,D+16z=0.$ (6.14)
[…]